# Benign Overfitting for $\alpha$ Sub-exponential Input

### v1

@@ Abstract
This paper investigates the phenomenon of benign overfitting in binary classification problems with heavy-tailed input distributions. We extend the analysis of maximum margin classifiers to $\alpha$ sub-exponential distributions, where $\alpha\in(0,2]$, generalizing previous work that focused on sub-gaussian inputs. Our main result provides generalization error bounds for linear classifiers trained using gradient descent on unregularized logistic loss in this heavy-tailed setting. We prove that under certain conditions on the dimensionality $p$ and feature vector magnitude $\|\mu\|$, the misclassification error of the maximum margin classifier asymptotically approaches the noise level. This work contributes to the understanding of benign overfitting in more robust distribution settings and demonstrates that the phenomenon persists even with heavier-tailed inputs than previously studied.
@@ 1 Introduction
Particularly in binary classification problems, a standard mixture model setting is often considered [6, 30, 11, 33]. This setting involves the classification of well-separated data under adversarially corrupted labels, where the input distribution is assumed to be sub-gaussian. However, benign overfitting in settings with input distributions heavier than sub-gaussian, that is, settings more robust to input variations, has not been extensively discussed.
Interestingly, numerical experiments suggest that benign overfitting can occur even in standard mixture settings with heavier-tailed distributions than the normal distribution, as illustrated in Figure 1. This observation further emphasizes the need to investigate benign overfitting in a broader range of distributional settings.
@@
Figure 1: Training and Test Error vs. Dimension ($p$) for Different $\gamma$ Values. The plot shows the test error and train error versus the dimension $p$ for a maximum margin classifier. The number of training samples $n_{\mathrm{train}}=200$ is kept constant, and the number of test samples $n_{\mathrm{test}}$ is $1000$. The number of parameters $p$ varies from $100$ to $1500$. The data is generated from a relaxed setting of a standard mixture model; two classes form a symmetric mixture of generalized normal distributions (see Example 2) with location parameter $x_{0}=[1,\cdots,1]$ and $x_{0}=[-1,\cdots,-1]$, shape parameter $\gamma=0.25,0.50,2$, controlling the tail heaviness, and scale parameter $\sigma$ adjusted so that the variance becomes 1 for each $\gamma$. The noise level $\eta$ is set to $0.05$, represented by the dotted line. The solid and dashed lines show the training and test errors respectively for different tail heaviness $0.25$ (blue), $0.5$ (green), and $2$ (red), with error bars indicating variability across $50$ trials. The training error remains near zero across all dimensions, while the test error initially decreases and then stabilizes around the noise level as the dimension increases.
@@
In this work, we consider binary classification where the input distribution has heavier tails than sub-gaussian, specifically $\alpha$ sub-exponential with $\alpha\in(0,2]$. The research aims to prove generalization error bounds demonstrating benign overfitting for a linear classifier using gradient descent on the unregularized logistic loss. This work generalizes the bound of Chatterji and Long [6].
@@ 1.1 Related works
Benign overfitting in classification: Most related to our work is the theoretical analysis of benign overfitting in classification settings. This line of research aims to understand why classifiers that perfectly fit noisy training data can still generalize well to unseen data. Wang and Thrampoulidis [31] studied a setting where the two classes are a symmetric mixture of gaussian (or sub-gaussian) distributions, without label noise. Chatterji and Long [6] studied overparameterized linear logistic regression on sub-gaussian mixture models with label flipping noise. They showed how gradient descent can train these models to achieve nearly optimal population risk. Cao et al. [5] extended this work, tightening the upper bound in the case without label flipping noise and establishing a matching lower bound for overparameterized maximum margin interpolators. Wang et al. [30] extended the analysis of maximum margin classifiers to multiclass classification in overparameterized settings. Frei et al. [11] proved that a two-layer fully connected neural network exhibits benign overfitting under certain conditions, such as well-separated log-concave distribution and smooth activation function. Cao et al. [4] focused on the benign overfitting of two-layer convolutional neural networks.
@@ 2.1 Notation
In this paper, we use the notation $[n]$ to denote the set $\{1,2,\ldots,n\}$ for a positive integer $n$. For a vector $x$, we use $\|x\|$ to denote its $\ell^{2}$ norm. For a matrix $A$, we use $\|A\|_{\mathrm{HS}}$ to denote its Hilbert–Schmidt norm and $\|A\|_{\mathrm{op}}$ to denote its operator norm.
@@ 2.2 $\alpha$ sub-exponential random variable
Definition 1 ($\alpha$ sub-exponential random variable).
@@ 2.3 Data generation process
    • $P_{\mathrm{clust}}:=P_{\mathrm{clust}}^{(1)}\times\cdots\times P_{\mathrm{clust}}^{(p)}$ is an arbitrary product distribution on $\mathbb{R}^{p}$ whose marginals are all mean-zero with the exponential Orlicz norm at most one, $\|X\|_{\psi_{\alpha}}$ if $X\sim P_{\rm clust}^{(j)}$.
    • For some $\kappa>0$, it holds that $\mathbb{E}_{q\sim P_{\mathrm{clust}}}[\|q\|^{2}]\geq\kappa p$.
@@
    Let $\tilde{P}$ be the distribution followed by $(\tilde{x},\tilde{y})$.
@@ 2.4 Maximum margin algorithm
$\theta^{(t+1)}=\theta^{(t)}-\beta\nabla R(\theta^{(t)})\quad\mathrm{where}\quad R(\theta):=\sum_{i=1}^{n}\log(1+\exp(-y_{i}\theta\cdot x_{i})),$
@@
They have proved this for a certain loss function with a certain smoothness.
@@ 2.5 Assumptions
We assume that $\alpha$ and $\kappa$ are fixed constants. We will prove that there exists large enough $C$, depending only on $\alpha$ and $\kappa$, such that our theorem holds for $C$.
We assume the following assumptions:
@@
    The dimension satisfies $p\geq C\max\{\|\mu\|^{2}n,n^{2}\left(\log\frac{n}{\delta}\right)^{\frac{2}{\alpha}}\}$,
@@
    where $c_{1}=\max\left(\frac{8}{\kappa},\frac{8}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\kappa+2\right).$
When $\alpha=2$, our assumptions correspond to that of Chatterji and Long [6].
@@ 3 Main Results
Building on these results, we can further analyze the implications of Theorem 4. This theorem reveals the relationship between the number of dimensions $p$ and $\|\mu\|$ in determining the success of learning. Specifically, when $\|\mu\|$ increases as $\|\mu\|=\Theta(p^{d})$ for any $d\in(1/4,1/2]$, the misclassification error of the maximum margin classifier asymptotically approaches the noise level $\eta$. This suggests that in high-dimensional feature spaces, if the signal is sufficiently strong, learning can be achieved while minimizing the impact of noise. These findings align with previous research. The rate of increase in $\|\mu\|$ for benign overfitting corresponds to that proved by Chatterji and Long[6] when $\alpha=2$.
@@
There exists a constant $c>0$ such that, under the assumptions (A1)-(A5), in the noisy rare-weak model, for any $\lambda\geq 0$ and all large enough $C$, with probability $1-\delta$, a maximum margin classifier $w$ satisfies
@@
Let $\lambda$ be a constant. Jin [14]demonstrated that for the noiseless rare-weak model, learning is impossible when $s=O(\sqrt{p})$ under the Gaussian assumption. Considering the fact that the gaussian distribution is an $\alpha$ sub-gaussian for every $\alpha$ in $(0,2]$, their counterexample can show that our upper bound has optimality in a sense. Strictly speaking, to fit Jin’s model to our model, we need to adjust the scale parameter $\sigma$ of $P_{\text{clust}}^{(j)}$ to make the exponential Orlicz norm less than 1. However, this adjustment does not affect the accuracy of the maximum margin classifier.
4 Proof of Theorem 4
In the Propositions and Lemmas of this section, we assume (A1)-(A5). For simplicity, we assume $U=I$. This assumption can be made without loss of generality for the following reasons:
@@ 4 Proof of Theorem 4
Let $\mathcal{S}:=\{(x_{1},y_{1}),\cdots,(x_{n},y_{n})\}$ be samples drawn according to $P$. We define the sets of indices of “noisy” and “clean” samples.
@@
Then, $\xi_{k}$ and $\tilde{\xi}_{k}$ are $\alpha$ sub-exponential, and the following proposition holds:
Proposition 8.
@@
    $\mathbb{E}[z_{k}]=\mathbb{E}[\tilde{z}_{k}]=\mu$.
@@
Proof.
By definition of $\tilde{z}_{k}$ and $z_{k}$,
Together with the definition of $\xi_{k}$,
@@
    $\underset{(\tilde{x},\tilde{y})\sim\tilde{P}}{\mathbb{E}}[\tilde{y}(w\cdot\tilde{x})]$, which equals $w\cdot\mu$.
@@
Proof.
Following the proof of Chatterji and Long [6], we have
$=\eta+\mathbb{P}_{(\tilde{x},\tilde{y})\sim\tilde{P}}\left[\left(\frac{w}{\|w\|}\cdot\tilde{\xi}\right)<-\frac{w}{\|w\|}\cdot\mu\right]$
Applying Proposition 16, there exists a positive constant $c$ such that
which completes our proof. ∎
@@
For any $\alpha\in(0,2]$ and $\kappa\in(0,1)$, there exists a constant $c_{1}\geq 1$ such that, for any $c^{\prime}$, for all large enough $C$, with probability at least $1-\delta$, the following hold:
@@
The proof of this lemma is in Appendix A.2.
From here on, we will assume that samples satisfy all the conditions of Lemma 10. Next, we will prove that the ratio of the losses between any pair of points is bounded. In this proof, we use Lemma 11.
Lemma 11.
For any $s_{1},s_{2}\in\mathbb{R}$,
First, we prove that the ratio of losses is bounded when the loss function is the exponential loss.
Lemma 12.
for all iterations $t\geq 0$,
where $c_{1}$ is a constant which satisfies Lemma 10.
Proof.
×$\quad\times\exp\left(-\beta\left(\sum_{k\neq 1}\frac{z_{k}\cdot z_{1}}{1+\exp(\theta^{(t)}\cdot z_{k})}-\sum_{k\neq 2}\frac{z_{k}\cdot z_{2}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)\right)$
$\frac{p}{c_{1}}\leq\|z_{k}\|^{2}\leq c_{1}p,$
Thus,
Now we consider two disjoint cases.
Case 1 ($A_{t}<2c_{1}^{2}$):
$\leq A_{t}\exp\left(\beta\left(c_{1}p+2nc_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)\right)$
Taking $\beta$ small enough that
we have
Case 2 ($A_{t}\geq 2c_{1}^{2}$):
By Lemma 11 and the induction hypothesis,
$\leq A_{t}\exp\left(-\frac{\beta c_{1}}{1+\exp(\theta^{(t)}\cdot z_{2})}\left(p-8c_{1}^{2}n\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)\right)$
By Assumption (A3) and (A4), for large enough $C$,
Thus,
$A_{t+1}<A_{t}\leq 4c_{1}^{2}.$
This completes the proof of the inductive step. ∎
Next, we prove the bound on the ratio of losses when the loss function is the sigmoid loss.
@@
Proof.
This lemma is derived from Lemma 11 and Lemma 12. ∎
Now, we prove the lower bound of $\mu\cdot w$.
@@
In the proof of Lemma 14, we use Lemma 15.
Lemma 15 (Soudry et al., 2018[27]).
For any linearly separable $\mathcal{S}$ and for all small enough learning rate $\beta$, we have
Proof of Lemma 14.
By Lemma 10,
By $|\mathcal{N}|\leq(\eta+c^{\prime})n$ and Lemma 13,
Thus, we have
By using this inequality repeatedly and $\theta^{(0)}=0$,
By taking the large-$t$ limit and using Lemma 15,
By definition of gradient descent iterations,
$\leq\beta\sum_{m=0}^{t}\|\nabla R(\theta^{(m)})\|$
With inequality (1), we have
$\mu\cdot w\geq\frac{\|w\|\|\mu\|^{2}}{4c_{1}\sqrt{p}},$
which completes our proof. ∎
By Lemma 9 and Lemma 14, we have Theorem 4.

### v2

# Benign Overfitting under Learning Rate Conditions for

@@ Abstract
This paper investigates the phenomenon of benign overfitting in binary classification problems with heavy-tailed input distributions, extending the analysis of maximum margin classifiers to $\alpha$ sub-exponential distributions ($\alpha\in(0,2]$). This generalizes previous work focused on sub-gaussian inputs. We provide generalization error bounds for linear classifiers trained using gradient descent on unregularized logistic loss in this heavy-tailed setting. Our results show that, under certain conditions on the dimensionality $p$ and the distance between the centers of the distributions, the misclassification error of the maximum margin classifier asymptotically approaches the noise level, the theoretical optimal value. Moreover, we derive an upper bound on the learning rate $\beta$ for benign overfitting to occur and show that as the tail heaviness of the input distribution $\alpha$ increases, the upper bound on the learning rate decreases. These results demonstrate that benign overfitting persists even in settings with heavier-tailed inputs than previously studied, contributing to a deeper understanding of the phenomenon in more realistic data environments.
@@ 1 Introduction
In the context of binary classification, a standard mixture model is often used to study benign overfitting [7, 12, 38]. This model involves classifying well-separated data with adversarially corrupted labels, assuming the input distribution is sub-gaussian. However, benign overfitting in settings with input distributions heavier than sub-gaussian, that is, settings more robust to input variations, has not been extensively discussed.
Figure 1: Boxplot of estimated tail index $\xi$ for feature vector components extracted from the intermediate layers of a CNN with ReLU activation, trained on various datasets (CIFAR-10 [17], CIFAR-100 [17], Fashion-MNIST [36], SVHN [28]). The tail index $\xi$ represents the heaviness of the distribution tails, with smaller values indicating heavier tails. The Gaussian and Exponential distributions are included for comparison purposes and were not passed through the CNN. The results indicate that the feature vectors for certain datasets, have heavier-tailed distributions than the Gaussian distribution. Further details are found in Appendix C.1.
@@
Figure 2: Training and test errors versus dimension $p$ for a maximum margin classifier. $n_{\mathrm{train}}=200$, $n_{\mathrm{test}}=1000$, $p$ ranges from 100 to 1500. Data is generated from a heavy-tailed setting using a generalized normal distribution, as detailed in Section 2.3 and Appendix C.2. The shape parameters are $\gamma=0.25,0.5,2$, with variance normalized to 1. Noise level $\eta$ is $0.05$ (dotted line). Solid and dashed lines show training and test errors, respectively, with $95$% confidence intervals as error bars over 50 trials. Training error remains near zero, while test error stabilizes around the noise level as $p$ increases.
Our numerical experiments indicate that feature vectors in convolutional neural networks (CNNs) with ReLU activation often exhibit distributions with tails heavier than sub-gaussian. The tail index $\xi$ can be intuitively understood as a parameter that characterizes the heaviness of the tails of a distribution. Specifically, for large values of $t$, the tail probability can be approximated as $\mathbb{P}[|X|>t]\simeq a\cdot\exp(-b\cdot t^{\xi})$, where smaller values of $\xi$ correspond to heavier tails in the distribution. Figure 1 indicates that these feature vectors have significantly heavy-tailed components. This finding emphasizes the necessity to extend benign overfitting analysis to accommodate a wider range of distributional settings. A more detailed explanation is provided in Appendix C.1.
Moreover, our numerical experiments suggest that benign overfitting can occur even in mixture model settings where the distributions have heavier tails than the normal distribution, as seen in Figure 2. Further details are available in Appendix C.2. This motivates the exploration of benign overfitting in more general distributional frameworks.
@@
In this work, we focus on binary classification tasks where the input distribution is $\alpha$ sub-exponential with $\alpha\in(0,2]$, implying tails heavier than sub-gaussian. We aim to establish generalization error bounds that demonstrate benign overfitting for a linear classifier trained using gradient descent on the unregularized logistic loss. Moreover, we derive an upper bound on the learning rate, a factor previously unexamined in this context, which plays a crucial role in demonstrating benign overfitting.
In this paper, we focus on the setting of Chatterji and Long [7], a pioneering work of benign overfitting theory on a simple model. Their results are extended to the heavy-tailed setting, and a more detailed discussion on the learning rate is provided.
@@ 1.1 Related works
Benign overfitting in classification: Most related to our work is the theoretical analysis of benign overfitting in classification settings. This line of research aims to understand why classifiers that perfectly fit noisy training data can still generalize well to unseen data. Wang and Thrampoulidis [35] studied a setting where the two classes are symmetric mixture of Gaussian (or sub-gaussian) distributions, without label noise. Chatterji and Long [7] studied overparameterized linear logistic regression on sub-gaussian mixture models with label flipping noise. They showed how gradient descent can train these models to achieve nearly optimal population risk. Cao et al. [6] extended this work, tightening the upper bound in the case without label flipping noise and establishing a matching lower bound for overparameterized maximum margin interpolators. Wang et al. [34] extended the analysis of maximum margin classifiers to multiclass classification in overparameterized settings. Frei et al. [12] proved that a two-layer fully connected neural network exhibits benign overfitting under certain conditions, such as well-separated log-concave distribution and smooth activation function. Cao et al. [5] focused on the benign overfitting of two-layer convolutional neural networks.
@@ 2.1 Notation
In this paper, we use the notation $[n]$ to denote the set $\{1,2,\ldots,n\}$ for a positive integer $n$. For a vector $x$, we use $\|x\|$ to denote its $\ell^{2}$ norm. For a matrix $A$, we use $\|A\|_{\mathrm{HS}}$ to denote its Hilbert–Schmidt norm and $\|A\|_{\mathrm{op}}$ to denote its operator norm. We use $s_{k}(A)$ to denote the $k$-th largest singular value of $A$. We use $O(\cdot)$ and $\Theta(\cdot)$ to refer to big-O and big-Theta notation.
@@ 2.2 $\alpha$ sub-exponential random variable
Definition 1 ($\alpha$ sub-exponential random variable, [29]).
@@ 2.3 Data generation process
    • $P_{\mathrm{clust}}:=P_{\mathrm{clust}}^{(1)}\times\cdots\times P_{\mathrm{clust}}^{(p)}$ is an arbitrary product distribution on $\mathbb{R}^{p}$ whose marginals are all mean-zero with the exponential Orlicz norm at most $1$, i.e., $\|X\|_{\psi_{\alpha}}\leq 1$ if $X\sim P_{\rm clust}^{(j)}$.
    • For some $\kappa>0$, it holds that
      $\mathbb{E}_{q\sim P_{\mathrm{clust}}}[\|q\|^{2}]\geq\kappa p.$
@@
    Let $\tilde{P}$ be the distribution of $(\tilde{x},\tilde{y})$.
@@ 2.4 Maximum margin algorithm
$\theta^{(t+1)}=\theta^{(t)}-\beta\nabla R(\theta^{(t)})$
$\quad\mathrm{where}\quad R(\theta):=\sum_{i=1}^{n}\log(1+\exp(-y_{i}\theta\cdot x_{i})),$
@@
They have proved this for a class of loss functions with certain smoothness.
@@ 2.5 Assumptions
We assume that $\alpha$ and $\kappa$ are fixed constants. Let $X=[y_{1}x_{1},\cdots,y_{n}x_{n}]$, where $\{(x_{k},y_{k})\}_{k=1}^{n}$ are samples drawn according from the distribution $P$. We will prove the main theorem and corollaries under the following assumptions with a sufficiently large constant $C$ depending only on $\alpha$ and $\kappa$.
@@
    The dimension satisfies
    $p\geq C\max\left(\|\mu\|^{2}n,n^{2}\left(\log\frac{n}{\delta}\right)^{\frac{2}{\alpha}}\right),$
@@
    $\beta\leq$ $\min\Bigg{(}8(s_{1}(X))^{-2},$
    $\frac{1}{c_{2}}\left(p+2n\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1}\Bigg{)},$
    where $c_{2}=2\max\left(\frac{8}{\kappa},\frac{8}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\kappa+2\right).$
When $\alpha=2$, assumptions (A1)-(A4) correspond to those in Chatterji and Long [7].
Due to assumption (A1), if we require a lower failure probability, $C$ must be set large, which in turn requires tighter lower bounds for $n$, $p$ and $\|\mu\|$ in assumptions (A2)-(A4). Moreover, as the tail heaviness of the distribution grows (i.e., as $\alpha$ decreases), the lower bounds for $n$, $p$, and $\|\mu\|$ become tighter, and the upper bound on the learning rate $\beta$ becomes more restrictive.
@@ 3.1 Generalization bound
The proof of this theorem is provided in Section 4. This theorem reveals the relationship between the number of dimensions $p$ and $\|\mu\|$ in determining the success of learning. Specifically, when $\|\mu\|$ increases as $\|\mu\|=\Theta(p^{\tau})$ for any $\tau\in(1/4,1/2]$, the misclassification error of the maximum margin classifier asymptotically approaches the noise level $\eta$. The rate of increase in $\|\mu\|$ for benign overfitting is same as that proved by Chatterji and Long [7] when $\alpha=2$. Therefore, our result shows that in high-dimensional feature spaces, if the signal is sufficiently strong, learning can be achieved while minimizing the impact of noise even for heavier tails ($\alpha<2$).
@@
There exists a constant $c>0$ such that, under assumptions (A1)-(A5), in the generalized noisy rare-weak model, for any $\lambda\geq 0$ and all large enough $C$, with probability $1-\delta$, a maximum margin classifier $w$ satisfies
@@
We will consider $\lambda$ as fixed. Jin [15] demonstrated that for the noiseless rare-weak model, learning is impossible when $s=O(\sqrt{p})$ under the Gaussian assumption. Considering the fact that the Gaussian distribution is an $\alpha$ sub-gaussian for every $\alpha$ in $(0,2]$, their counterexample can show that our upper bound has optimality in a sense. Strictly speaking, to fit Jin’s model to our model, we need to adjust the scale parameter $\sigma$ of $P_{\text{clust}}^{(j)}$ to make the exponential Orlicz norm less than 1. However, this adjustment does not affect the accuracy of the maximum margin classifier.
3.2 Learning rate
We perform a detailed analysis of sufficient conditions for the learning rate when benign overfitting occurs. To concretely calculate the assumption of the learning rate, a bound of the largest singular value of $X$ is used.
Proposition 6 (A bound of the singular values of $X$).
For any $\delta\geq 0$, with probability at least $1-\delta$, there are constants $c_{5}$, $c_{6}$, $c_{7}$, $c_{8}$ depending only on $\alpha$ such that
$s_{1}(X)$
$\leq\sqrt{p}\Bigg{(}c_{5}+\frac{c_{6}\sqrt{n}}{p}\sum_{i=1}^{p}|\mu_{i}|+\frac{2n\|\mu\|^{2}}{p}$
$\quad+\frac{c_{7}+c_{8}\max_{i}|\mu_{i}|\sqrt{n}}{p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}}\Bigg{)}.$
The proof of Proposition 6 is in Appendix B.2. According to Proposition 6, a sufficient condition for assumption (A5) can be expressed as assumption (A6):
  1. (A6)
    The learning rate satisfies:
    $\beta\leq$
    $\min\Bigg{(}\frac{8}{p}\Bigg{(}c_{5}+\frac{c_{6}\sqrt{n}}{p}\sum_{i=1}^{p}|\mu_{i}|+\frac{2n\|\mu\|^{2}}{p}$
    $\quad+\frac{c_{7}+c_{8}\max_{i}|\mu_{i}|\sqrt{n}}{p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}}\Bigg{)}^{-2},$
    $\quad\frac{1}{c_{2}p}\left(1+\frac{2n}{p}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1}\Bigg{)}.$
By using assumption (A6) instead of (A5), we obtain Corollary 7.
Corollary 7.
Under assumptions (A1)-(A4) and (A6) for all large enough $C$, with probability at least $1-2\delta$, the same generalization error bound as in Theorem 4 holds.
Moreover, by Corollary 7, we obtain Corollary 8 and 9. The proofs of Corollaries 8 and 9 are in Appendix B.3.
Corollary 8.
Under assumptions (A1)-(A4) for all large enough $C$, if $\beta$ satisfies
$\beta\leq c_{9}p^{-1}$
where $c_{9}$ is a constant depending on $\alpha,\kappa,\delta$, and $n$, with probability at least $1-2\delta$, the same generalization error bound as in Theorem 4 holds.
This corollary implies that when $p$ and $\|\mu\|$ grow large while $n$ and $\delta$ are fixed under assumptions (A3) and (A4), $\beta=O(p^{-1})$ is sufficient for the same result as Theorem 4. The order remains unchanged even when $\alpha$ is small.
Corollary 9.
Under assumptions (A1)-(A4) for all large enough $C$, if $\beta$ satisfies
$\beta\leq c_{10}p^{-1}\left(1+n^{\frac{2}{\alpha}-1}(\log n)^{-\frac{1}{\alpha}}\right)^{-2}$
where $c_{10}$ is a constant depending on $\alpha,\kappa$, and $\delta$, with probability at least $1-2\delta$, the same generalization error bound as in Theorem 4 holds.
Since $p\geq n^{2}$, it is straightforward to show that $\beta=O(p^{-\frac{2}{\alpha}})$ ensures the same generalization error bound as not only $p$ and $\|\mu\|$, but also $n$ grows under assumptions (A3) and (A4). As $\alpha$ decreases, the order also decreases, indicating that, for heavy-tailed distributions, the learning rate must be reduced accordingly.
4 Sketch of proof of Theorem 4
In the lemmas of this section, we assume (A1)-(A5). The proofs of the lemmas in this section are provided in Appendix B.1. For simplicity, we assume $U=I$. This assumption can be made without loss of generality for the following reasons:
@@ 4 Sketch of proof of Theorem 4
We define the sets of indices of “noisy” and “clean” samples.
@@
Then, $\xi_{k}$ and $\tilde{\xi}_{k}$ are $\alpha$ sub-exponential, and the following lemma holds:
Lemma 12.
@@
    $\mathbb{E}[z_{k}]=\mathbb{E}[\tilde{z}_{k}]=\mu$ and
@@
    $\underset{(\tilde{x},\tilde{y})\sim\tilde{P}}{\mathbb{E}}[\tilde{y}(w\cdot\tilde{x})],$
    which equals $w\cdot\mu$.
@@
For any $\alpha\in(0,2]$ and $\kappa\in(0,1)$, there exists a constant $c_{1}\geq 1$ such that, for any $c^{\prime}$, for all large enough $C$, with probability at least $1-\delta$, the following holds:
@@
From here on, we will assume that samples satisfy all the conditions of Lemma 14.
The next lemma provides the bound on the ratio of losses when the loss function is the sigmoid loss.
@@
Soudry et al. [30] provide results regarding the convergence behavior of $\theta^{(t)}$ when the data is linearly separable.
Lemma 16 (Soudry et al., 2018 [30]).
For any linearly separable $\mathcal{S}$ and for $\beta\leq 8(s_{1}(X))^{-2}$, we have
Using Lemmas 14, 15, and 16, we derive Lemma 17.
@@
By Lemmas 13 and 17, we have Theorem 4.
5 Simulation
We conducted simulation studies to assess the performance of the maximum margin classifier across various conditions, specifically focusing on how dimensionality, tail heaviness, and learning rates interact. The simulation was designed with the following parameters: the training set consisted of $500$ samples, we used $1000$ test samples to assess generalization, and each experiment was repeated $5$ times, and the results were averaged to ensure robustness. The link to the detailed code for the experiments is provided in Appendix D.
Figure 3: A heatmap showing the mean test error for $\beta=0.001$ with the horizontal axis representing the dimension $p$ and the vertical axis representing the shape parameter $\gamma$.
Figure 4: A heatmap showing the mean test error for $\gamma=0.8$ with the horizontal axis representing the dimension $p$ and the vertical axis representing the learning rate $\beta$.
Figure 5: A heatmap showing the mean test error for $p=4000$ with the horizontal axis representing the shape parameter $\gamma$ and the vertical axis representing the learning rate $\beta$.
5.1 Data generation
The data was generated by the heavy-tailed setting, as described in Section 2.3. We set $P_{\mathrm{clust}}$ to be the generalized normal distribution with a scale parameter of $1$ and a shape parameter $\gamma$ ranging from $0.5$ to $1$. This distribution is used to control the tail behavior of the data, where smaller values of $\gamma$ correspond to heavier tails.
For the mean vector, $\mu$, the first $\lfloor p^{2/3}\rfloor$ elements were set to 1, and the remaining elements were set to 0, ensuring that $\|\mu\|=\Theta(p^{1/3})$. We chose an orthogonal matrix $U$ such that $U=I$, the identity matrix.
We also incorporated label noise by flipping the labels with a noise level of $\eta=0.05$, meaning that each true label was flipped with a probability of $\eta$.
5.2 Model training
We used the maximum margin classifier, as described in Section 2.4. The model was trained for $100000$ epochs to ensure convergence. We conducted three different numerical experiments to observe how these conditions influence test error.
5.3 Results and discussion
Experiment 1: interaction between dimension $p$ and shape parameter $\gamma$ (Figure 3)
In the first experiment, we investigated how the interaction between the dimensionality $p$ and the shape parameter $\gamma$ influences model performance. As $p$ increases, the test error initially decreases and then stabilizes around the noise level. For smaller values of $\gamma$ (heavier tails), the stabilization occurs more slowly, indicating that learning from heavy-tailed distributions is more challenging. In contrast, for larger values of $\gamma$ (lighter tails), the model converges faster.
These results suggest that high-dimensional parameter spaces permit benign overfitting, regardless of the tail heaviness. However, for heavier-tailed distributions (smaller $\gamma$), more dimensions are required to achieve similar performance compared to lighter-tailed distributions. This is consistent with the theoretical assumptions (A3) and (A4).
Experiment 2: impact of dimension $p$ and learning rate $\beta$ (Figure 4)
In the second experiment, we explored how the interaction between dimensionality $p$ and learning rate $\beta$ affects the test error. For large $p$, benign overfitting does not occur unless a small learning rate $\beta$ is chosen. If $\beta$ is too large, the learning process struggles to make progress.
As $p$ increases, the generalization error should decrease, as predicted by the benign overfitting bound. However, if the learning rate $\beta$ is not sufficiently small, the conditions outlined in Corollary 7 are not satisfied, and the learning process does not perform well. Our simulations suggest that in high-dimensional parameter spaces, the learning rate $\beta$ must be reduced to enable benign overfitting.
Experiment 3: impact of shape parameter $\gamma$ and learning rate $\beta$ (Figure 5)
In the third experiment, we fixed the dimensionality at $p=4000$ and examined the interaction between the shape parameter $\gamma$ and the learning rate $\beta$. For smaller $\gamma$ (heavier tails), the model is more sensitive to the choice of $\beta$. In particular, larger values of $\beta$ result in higher test errors for smaller $\gamma$. Conversely, for larger values of $\gamma$ (lighter tails), the model performs well even with larger learning rates. This suggests that careful tuning of the learning rate is crucial when dealing with heavy-tailed distributions to achieve benign overfitting.
These findings align with the theoretical results, indicating that when $\gamma$ is small, if $\beta$ is not sufficiently small, the condition on $\beta$ specified in Corollary 9 is violated.